OpenStudy (anonymous):

1 question help! medal

OpenStudy (anonymous):

Yes?

OpenStudy (anonymous):

For each function, what is the output of the given input? For f(x)= 4x+7 , find f(4)

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

23

OpenStudy (anonymous):

ty @grray7 ! cn you show me how you got 23?

OpenStudy (anonymous):

well put f(4) in for f(x) and then another 4 where you see x in the equation, so, 4*4=16 +7=23

OpenStudy (anonymous):

ty :)

OpenStudy (anonymous):

yep